OpenStudy (anonymous):

Find the inverse of the function. f(x) = x^3 + 4

OpenStudy (xapproachesinfinity):

set y=x^3+4 switch x and y role x=y^3+4 and solve for y

OpenStudy (anonymous):

Why do you substitute y for f(x)?

OpenStudy (anonymous):

@xapproachesinfinity i got y=square root of -4 + x

OpenStudy (xapproachesinfinity):

No you should get \(y=\sqrt[3]{x-4}\) so the \(f^{-1}(x)=\sqrt[3]{x-4}\) cube root not square root

OpenStudy (anonymous):

ohhhhh okay thanks!

OpenStudy (xapproachesinfinity):

Welcome^_^ Pay attention to Radicals -_^